OpenStudy (anonymous):

The product of two consecutive odd integers is 195. Find the integers.

OpenStudy (lgbasallote):

hmm..well let's say the first integer is x so the next consecutive ODD integer would be x+2 yes? so... x(x+2) = 195 since the product of those two is 195 do you know what to do next?

OpenStudy (anonymous):

x^2+2=195?

OpenStudy (anonymous):

hell i would guess and check!

OpenStudy (inkyvoyd):

Find a rough estimate of the square root of 200.

OpenStudy (anonymous):

99 + 101 = 200 too big 95 + 97 = 192 too small try one more and i bet you get it

OpenStudy (inkyvoyd):

Then, guess and check. This isn't the way you are supposed to, but it is much faster.

OpenStudy (inkyvoyd):

sqrt(200) is about 14. So, try 13*15

OpenStudy (inkyvoyd):

Which is your answer.

OpenStudy (anonymous):

oh product, not sum ok how about \(13\times 15\)

OpenStudy (inkyvoyd):

because, (14-1)(14+1)=14^2-1^1=196-1=195

OpenStudy (anonymous):

wow got it on the first try

OpenStudy (anonymous):

Lol thanks everyone

OpenStudy (lgbasallote):

lol @satellite73 =))))) nice strategy :P though it's product so it gets big...it gets slower hahaha

OpenStudy (inkyvoyd):

No, actually, two consectutive odd numbers is always easy

OpenStudy (inkyvoyd):

You can view it as (x)(x+2) where x is the smaller number OR you can view it as (x-1)(x+1), where x is the even number in between.

OpenStudy (inkyvoyd):

then, it's x^2-1=a or, x=plusminus sqrt(1+a)

OpenStudy (anonymous):

well lets imagine that you want to solve an equation, and you write \(x(x+2)=195\) the next thing you do is write \(x^2+2x-195=0\) and then you think "can i find two numbers whose product is 195 that are two apart, which is exactly what the original problem asks! so there is nothing gained by writing an equation

OpenStudy (inkyvoyd):

satellite, from that equation, all you have to do is complete the square; add 196 to both sides.

OpenStudy (inkyvoyd):

whenever I see x^2+2ax..., I always just add a^2 and go with it.

OpenStudy (inkyvoyd):

but seriously, note the difference between x(x+2) and (x+1)(x-1) A little bit of thinking goes a long way...

OpenStudy (anonymous):

by complete the square i guess you could write \[x^2+2x=195\] \[(x+1)^2=196\] and then pray for a pefect square

OpenStudy (anonymous):

as luck would have it you get 14, but really i would for real guess and check. if you don't get it on the first try you are sure to get it on the second

OpenStudy (inkyvoyd):

Ok, how bout the product of an 2 integers, both odd, with one being 6 greater than the other? say the product is also 196...

OpenStudy (inkyvoyd):

*not 196, 187